OpenStudy (anonymous):

write the equation fo the line through(6,4) that is perpendicular to y = 1/3x - 2?

OpenStudy (amistre64):

what is an important property of perp lines?

OpenStudy (anonymous):

y = -3x + 22

OpenStudy (anonymous):

The clue here will be the line will have an equation y = mx + c, where m is the negative reciprocal of the line y =1/3x -2. then just plug in the values of x= 6 and y=4 to obtain c.

OpenStudy (anonymous):

i m getting y = 1/3x +2?

OpenStudy (lalaly):

y-4=-3(x-6) y-4=-3x+18 y=-3x+22

OpenStudy (amistre64):

y = mx -mPx +Py

OpenStudy (amistre64):

since perp lines have flipped slopes ... and swapped signs; 1/3 becomes -3 y = -3x -(-3)(6) + 4

OpenStudy (anonymous):

thx